Find a rational no between -3/7 and 1/2
1
Expert's answer
2019-07-31T09:55:22-0400

A number is rational when it can be represented as

"\frac p q,"

where p, q are integers. For example, 0 is a rational number because it can be represented as

"\frac 0 1"


We can say

"\frac c d \le \frac a b \le \frac e f"

only if

"\begin{cases}
 bc \le ad \\
 af \le be
\end{cases}"

where b, d and f are positive. Otherwise, the denominator and numerator of fractions with negative denominators should be multiplied by -1, thus the denominator becomes positive. For example,

"\frac {-3} 7 \le \frac 0 1 \le \frac 1 2"

is true because

"\begin{cases}
 -3 \le 0 \\
 0 \le 1
\end{cases}"